OpenStudy (idku):

A mass m1 = 4.8 kg rests on a frictionless table and connected by a massless string over a massless pulley to another mass m2 = 4.5 kg which hangs freely from the string. When released, the hanging mass falls a distance d = 0.77 m.

OpenStudy (idku):

I can't figure out: (Part 3) What is the final speed of the two blocks? (Part 4) How much work is done by tension on m1?

OpenStudy (idku):

\(\color{#000000}{ \displaystyle F=ma }\) \(\color{#000000}{ \displaystyle F(\Delta t)^2=m\Delta x }\) \(\color{#000000}{ \displaystyle F(\Delta t)^2=(m_1+m_2)d }\) the force here is the weight of m2, \(\color{#000000}{ \displaystyle (m_2g)(\Delta t)^2=(m_1+m_2)d }\) \(\color{#000000}{ \displaystyle (9.80*4.5)(\Delta t)^2=(4.8+4.5)*0.77 }\) \(\Delta t=0.280360\) \(\color{#000000}{ \displaystyle F=ma }\) \(\color{#000000}{ \displaystyle F(\Delta t)=(\Delta v)m }\) \(\color{#000000}{ \displaystyle (9.80*4.5)*(0.402965)=(v_f-0)(4.8+4.5) }\) \(\color{#000000}{ \displaystyle v_f=1.91083 }\)

OpenStudy (idku):

this is what I attempted but turned out to be incorrect.